Circular No. 7438

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cm IN ANONYMOUS GALAXY S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that an image taken by A. Mossman on June 5.15 UT with the F. L. Whipple Observatory 1.2-m telescope (+ 4shooter) yields the following precise position for SN 2000cm (cf. IAUC 7436, 7437), α = 12^{h}12^{m}59^{s}.43, δ = +7^{o}18′00′′.6 (equinox 2000.0, uncertainty ± 0′′.3). H. Yamaoka, Kyushu University, reports the following position end figures, measured from an unfiltered CCD image taken on June 4 UT by F. M. Ewalt (East Killingly, CT) with a 0.25-m telescope: 59^{s}.43, 00′′.8.

M. Turatto, A. Pastorello, E. Cappellaro, and F. Cedrati, Osservatorio Astronomico di Padova, write: “We obtained spectra (range 350–750 nm, resolution 2 nm) on June 4.9 UT with the Asiago 1.82-m telescope (+ AFOSC), showing SN 2000cm to be a type-Ia supernova ∼ 2 weeks after maximum. The strongest features were emissions measured at 464, 539, and 592 nm. Based on comparison with spectra of other type-Ia events, we estimate the redshift of the parent galaxy to be ∼ 2500 km/s. The typical absorption of Si II is still visible and measured at 618 nm, corresponding to an expansion velocity of ∼ 10 000 km/s.”

SUPERNOVA 1998fe IN NGC 6027D W. D. Li, University of California at Berkeley, reports the discovery

with the 0.8-m Katzman Automatic Imaging Telescope of an apparent supernova in Lick Observatory Supernova Search (cf. IAUC 6627, 7126) archive images of NGC 6027D, which is one of the ‘Seyfert Sextet’ galaxies. SN 1998fe was present on unfiltered images taken on July 19.3, 21.3, 23.3, and 26.3 UT, all with mag ∼ 18.0, and it was located at α = 15^{h}59^{m}12^{s}.84, δ = +20^{o}45′39′′.0 (equinox 2000.0), which is 0′′.7 west and 2′′.4 north of the nucleus of NGC 6027D. SN 1998fe is not visible in subsequent images of the field taken in 1999 and 2000 (most of which have limiting mag ∼ 19.0).

SUPERNOVA 2000cj IN NGC 6753 R. H. McNaught, Australian National University, reports an improved

position for SN 2000cj (cf. IAUC 7425) from four CCD images taken with the Siding Spring 1.0-m telescope on June 12.64 UT (using 139 USNO CCD Astrograph Catalog stars, rms 0′′.14 in α and 0′′.15 in δ): α = 19^{h}11^{m}27^{s}.658, δ = −57^{o}03′19′′.41 (equinox J2000.0), which is 33′′.2 east and 21′′.8 south of the center of NGC 6753. The magnitude was R ∼ 15.7.

2000 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green